OpenStudy (anonymous):

Twice a number plus four equals 6 what is the number!

OpenStudy (niel):

1

OpenStudy (niel):

1x2+4=6

OpenStudy (niel):

the math is this: 2x+4=6 -4 -4 2x =2 ___________ 2 x = 1

OpenStudy (anonymous):

let the number be called Q (cause I like that letter) 2 x Q + 4 = 6 (move four over, so now youre subtracting) 2 x Q = 6 - 4 2 x Q = 2 (move two over so youre dividing) Q = 2/2 Q = 1 The number is one! Yay!